# A Survey on Self-play Methods in Reinforcement Learning

### v3

@@ 4.2.2 DouDiZhu
4.2.3 Mahjong
Mahjong has evolved into various global variants, including the famous Japanese version known as Riichi Mahjong. This game is typically played by four players who must navigate both the visible aspects of the game, such as discarded tiles, and the hidden elements, like their own hand and the unseen public tiles. The strategic depth and complexity of Mahjong pose significant challenges. Despite ongoing research, these findings have yet to reach expert human levels (Kurita & Hoki, 2020; Gao et al., 2019; Mizukami & Tsuruoka, 2015). Difficulties are navigating incomplete information, dynamically adapting strategies to multiple opponents, and contending with complex winning rules and an enormous number of possible game states, estimated at $10^{169}$ (Zha et al., 2019). Suphx (Li et al., 2020) is recognized as one of the first algorithms to master Mahjong, achieving a performance level comparable to expert human players, precisely 10 dan on Tenhou (Tsunoda, ), the most popular online Mahjong platform. Initially, Suphx employs supervised learning, utilizing expert data to train its model. It then advances its capabilities through vanilla self-play combined with RL. Similarly, NAGA (Village, ), developed by Dwango Media Village, and LuckyJ (Tencent, ), designed by Tencent, have also achieved the rank of 10 dan on Tenhou. Furthermore, LuckyJ has even defeated human professional players.
4.3 Video Games
@@
4.3.1 StarCraft II
StarCraft is a real-time strategy (RTS) game. It has three distinct species: the Terrans, Zerg, and Protoss, each with unique units and strategic options that enhance the gameplay complexity. Renowned for its balanced gameplay, strategic depth, and intense competitiveness, the game challenges players to gather resources, construct bases, and build armies. Victory requires meticulous planning and tactical execution, with defeat occurring when a player loses all their buildings. AlphaStar (Vinyals et al., 2019) dominates the 1v1 mode competitions in StarCraft II and has defeated professional players. Its framework is similar to AlphaGo (Silver et al., 2016), initially utilizing supervised learning to train the policy with expert data. Subsequently, it uses a hierarchical self-play method combined with end-to-end RL to train the networks. More specifically, The proposed self-play method divides all the agents into three types: main agents, league exploiters, and main exploiters. It maintains a policy pool of past players that records all these types of agents. Main agents engage in traditional self-play algorithms like FSP and PFSP, competing against main agents themselves and other agents in the policy pool. They are periodically added to the pool and never reset. League exploiters also use a traditional self-play method PFSP to play against all policy pool agents, added to the pool if they show a high win rate and potentially reset to expose global blind spots. Main exploiters only compete with main agents to improve their robustness, are added to the pool after achieving a high win rate or specific training steps, and are reset upon each addition. Among those three agent types, the main agent is the core agent and embodies the final AlphaStar strategy. However, the introduction of three agent types significantly increases the training computation. Further studies (Han et al., 2020; Wang et al., 2021; Huang et al., 2024) have enhanced the league self-play training procedure.
4.3.2 MOBA Games
@@
OpenAI Five (Berner et al., 2019) defeated the world champion team in a simplified version of Dota 2 that featured a limited pool of heroes and certain banned items. The training process introduces a self-play method that combines two traditional self-play algorithms: with an 80% probability of engaging in vanilla self-play and a 20% probability of employing a technique similar to the PFSP used in AlphaStar (Vinyals et al., 2019). This technique selects each policy from the policy pool based on its quality score, which is continuously updated from competition results throughout training. Higher quality scores increase the likelihood of a policy being selected. OpenAI Five also requires extensive training resources.
Another notable MOBA game, especially viral in China, is Honor of Kings. The 1v1 mode is conquered by Ye et al. (2020b), which boasts a significant winning rate against top professional players. It utilizes a traditional self-play method $\delta$-uniform self-play to train the RL networks. The 5v5 mode was later mastered by Ye et al. (2020a) as well. Unlike OpenAI Five, this work expands the hero pool to 40 heroes, exponentially increasing the possible combinations of lineups. It proposes a new self-play method referred to as curriculum self-play learning (CSPL). Specifically, the training process is divided into three stages. The first stage involves training fixed lineups through self-play and utilizing human data to balance the two teams to aid policy improvements. The second stage employs multi-teacher policy distillation to produce a distilled model. The final stage uses this distilled model as the initial model for another round of self-play with randomly picked lineups. This approach defeats professional player teams. The self-play generated data is also used to learn compelling lineup drafting by utilizing MCTS and neural networks (Chen et al., 2021).
4.3.3 Google Research Football
@@
WeKick (Ziyang Li, 2020) claimed victory in the GRF competition on Kaggle (Google, 2020), which simplifies the game dynamics by allowing competitors to control only one player, either the ball carrier on offense or the nearest defender on defense. It employed self-play strategies similar to those used in league training (Vinyals et al., 2019). It initializes its opponent policy pool using strategies developed through RL and Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon, 2016) to facilitate the training process.
@@
Further research delves into the full football game rather than the simplified version. Team-PSRO (McAleer et al., 2023) extends PSRO series of self-play algorithms to team games, outperforming baselines in the 4v4 version of the full GRF. In the context of the 11v11 version, where the goalkeeper is rule-based controlled, TiKick (Huang et al., 2021a) utilizes vanilla self-play to collect samples and then employs imitation learning. Fictitious Cross-Play (FXP) (Xu et al., 2023) proposes a new self-play method similar to AlphaStar (Vinyals et al., 2019). It introduces two populations: the main population and the counter population. Policies in the counter population improve solely by cross-playing with policies in the main population as opponents, while policies in the main population engage in playing with policies from both populations. FXP achieves a win rate of over 94% against TiKick. TiZero (Lin et al., 2023), a follow-up to TiKick, combines curriculum learning with FSP and PFSP (Vinyals et al., 2019) to avoid expert data reliance, achieving a higher TrueSkill rating (Herbrich et al., 2006) than TiKick.
Table 2: Overview of representative studies in empirical analysis.
Category Game Number of Perfect Number of Algorithms Self-play Method Search Expert Data
Players Information Game States
Board Game Chess 2 $\checkmark$ $10^{45}$ AlphaZero Vanilla SP $\checkmark$ $\times$
MuZero Vanilla SP $\checkmark$ $\times$
Go 2 $\checkmark$ $10^{360}$ AlphaGo FSP $\checkmark$ $\checkmark$
AlphaGo Zero Vanilla SP $\checkmark$ $\times$
AlphaZero Vanilla SP $\checkmark$ $\times$
MuZero Vanilla SP $\checkmark$ $\times$
Stratego 2 $\times$ $10^{535}$ DeepNash R-NaD $\times$ $\times$
Card Game HULHE 2 $\times$ $10^{17}$ Cepheus CFR+ $\checkmark$ $\times$
NFSP NFSP $\times$ $\times$
HUNL 2 $\times$ $10^{164}$ DeepStack CFR+ $\checkmark$ $\times$
Libratus MCCFR $\checkmark$ $\times$
ReBel CFR-D / FSP $\checkmark$ $\times$
Six-player No-limit Texas Hold’em 6 $\times$ $>10^{164}$ Pluribus MCCFR $\checkmark$ $\times$
DouDiZhu 3 $\times$ $10^{76}\sim 10^{106}$ DeltaDou Vanilla SP $\checkmark$ $\times$
DouZero Vanilla SP $\times$ $\checkmark$
PerfectDou Vanilla SP $\times$ $\times$
Mahjong 4 $\times$ $10^{169}$ Suphx Vanilla SP $\checkmark$ $\checkmark$
Video Game StarCraft II 2 $\times$ / AlphaStar FSP & PFSP $\times$ $\checkmark$
Dota 2 10 $\times$ OpenAI Five Vanilla SP & PFSP $\times$ $\times$
Honor of Kings 10 $\times$ MOBA AI $\delta$-uniform SP $\checkmark$ $\checkmark$
Google Research Football 22 $\checkmark$ TiZero FSP & PFSP $\times$ $\times$
5 Open Problems and Future Work
@@
Self-play methods have shown remarkable performance by leveraging iterative learning and adapting to complex environments. However, significant challenges and open questions remain, presenting opportunities for further research and development.
5.1 Theoretical Foundation
@@
Although NE has been shown to exist in games with finite players and finite actions (Nash et al., 1950), computing NE with self-play algorithms in larger games remains challenging and consequently, many studies aim to achieve approximate NE (Li et al., 2024). However, in some cases, even computing an approximate NE is difficult (Daskalakis, 2013). Some research has resorted to higher levels of equilibrium, such as CE (Marris et al., 2021) and $\alpha$-rank (Muller et al., 2020). Although many algorithms are developed with theoretical foundations, there is often a gap in complex real-world scenarios. For example, although self-play algorithms like AlphaGo (Silver et al., 2016), AlphaStar (Samvelyan et al., 2019), and OpenAI Five (Berner et al., 2019) achieve empirical success, they lack formal game theory proofs behind their effectiveness. Future work should focus on bridging this gap by developing new theoretically sound and practically effective algorithms or proving the theoretical underpinnings of existing successful self-play algorithms in complex environments.
5.2 Non-stationarity of the Environment
@@
5.3 Scalability and Training Efficiency
@@
The scalability of self-play methods faces significant challenges as the number of teams and players within those teams increases. As the number of participants grows, the complexity of interactions explodes. For example, in OpenAI Five (Berner et al., 2019), the hero pool size is limited to only 17 heroes. MOBA AI (Ye et al., 2020a) extends this to a 40-hero pool with the help of curriculum learning, but it still cannot cover the entire hero pool available in the actual game. One potential solution is leveraging players’ inherent connections to optimize the learning process. For instance, using graph-based models to represent and exploit the relationships between players can help manage and reduce the complexity of large-scale multi-agent environments. These scalability issues are fundamentally rooted in the limited training efficiency of self-play methods from two aspects including computation and storage. The first issue is computational efficiency, which is induced by the iterative nature of self-play where agents repeatedly play against themselves or past versions. Moreover, although forming more complex populations and competitive mechanisms (Samvelyan et al., 2019) can enhance the intensity and quality of training, it further exacerbates the demand for computational resources. Techniques such as parallel computing, distributed learning, and more efficient neural network architectures could be explored to address these challenges. The second issue is storage because self-play requires maintaining a policy pool. Even when using a shared network architecture, storing the parameters of large models can be problematic. This issue is particularly pronounced in regret-minimization-based self-play algorithms, which must store the regrets for each information set and potential action. Managing both the computational load and storage requirements is essential for improving the overall training efficiency and scalability of self-play methods.
5.4 With Large Language Models
@@
With their remarkable capability and emergent generalizability, large language models (LLMs) have been regarded as a potential foundation for achieving human-level intelligence (Achiam et al., 2023), and self-play methods have been proposed to fine-tune LLMs, enhance LLMs’ reasoning performance, and build LLM-based agents with strong decision-making abilities. Post-training fine-tuning (Ouyang et al., 2022; Bai et al., 2022) is a key step in aligning LLM with more desired behaviors, but it requires a huge amount of human-annotated data. To reduce reliance on human-annotated data, Self-play fIne-tuNing (SPIN) (Chen et al., 2024) introduces a self-play mechanism to generate training data using the LLM itself and fine-tuning the LLM by distinguishing self-generated responses from human-annotated data. Another line of work (Munos et al., 2024; Swamy et al., 2024; Wu et al., 2024) formulate the alignment problem as a two-player constant-sum game and use self-play methods to solve the game. Some other work (Huang et al., 2022; Yuan et al., 2024) also utilizes model-generated data to fine-tune LLMs with minimal human annotations. The idea of self-improvement has also been applied to improve the reasoning ability of LLMs. Self-Play of Adversarial Game (SPAG) (Cheng et al., 2025) observes that self-play in a two-player adversarial language game called Adversarial Taboo can boost the LLM’s performance on various reasoning benchmarks. Besides improving the capability of LLMs, self-play methods have also contributed to building LLM-based agents with strong strategic abilities. A representative work is Cicero († et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR) which achieves human-level play in the game of Diplomacy by combining language models with an RL policy trained by self-play. Cicero uses the self-play policy to produce an intended action and prompts the language model to generate languages conditioned on the policy’s intent. Another work (Xu et al., 2024) also combines LLM with self-play policy but takes a different approach by first prompting the LLM to propose multiple action candidates and then using the self-play policy to produce the action distribution over these candidates. Despite recent progress, applying self-play with LLMs is still underexplored and requires further research.
5.5 Real-World Applications

### v4

@@ 4.2.2. DouDiZhu
4.2.3. Mahjong
Mahjong has evolved into various global variants, including the famous Japanese version known as Riichi Mahjong. This game is typically played by four players who must navigate both the visible aspects of the game, such as discarded tiles, and the hidden elements, like their own hand and the unseen public tiles. The strategic depth and complexity of Mahjong pose significant challenges. Despite ongoing research, these findings have yet to reach expert human levels (Kurita and Hoki, 2020; Gao et al., 2019; Mizukami and Tsuruoka, 2015). Difficulties are navigating incomplete information, dynamically adapting strategies to multiple opponents, and contending with complex winning rules and an enormous number of possible game states, estimated at $10^{169}$ (Zha et al., 2019). Suphx (Li et al., 2020b) is recognized as one of the first algorithms to master Mahjong, achieving a performance level comparable to expert human players, precisely 10 dan on Tenhou (Tsunoda, [n. d.]), the most popular online Mahjong platform. Initially, Suphx employs supervised learning, utilizing expert data to train its model. It then advances its capabilities through vanilla self-play combined with RL. Similarly, NAGA (Village, [n. d.]), developed by Dwango Media Village, and LuckyJ (Tencent, [n. d.]), designed by Tencent, have also achieved the rank of 10 dan on Tenhou. Furthermore, LuckyJ has even defeated human professional players.
4.3. Video Games
@@
4.3.1. StarCraft II
StarCraft is a real-time strategy (RTS) game. It has three distinct species: the Terrans, Zerg, and Protoss, each with unique units and strategic options that enhance the gameplay complexity. Renowned for its balanced gameplay, strategic depth, and intense competitiveness, the game challenges players to gather resources, construct bases, and build armies. Victory requires meticulous planning and tactical execution, with defeat occurring when a player loses all their buildings. AlphaStar (Vinyals et al., 2019) dominates the 1v1 mode competitions in StarCraft II and has defeated professional players. Its framework is similar to AlphaGo (Silver et al., 2016), initially utilizing supervised learning to train the policy with expert data. Subsequently, it uses a hierarchical self-play method combined with end-to-end RL to train the networks. More specifically, The proposed self-play method divides all the agents into three types: main agents, league exploiters, and main exploiters. It maintains a policy pool of past players that records all these types of agents. Main agents engage in traditional self-play algorithms like FP and PFSP, competing against main agents themselves and other agents in the policy pool. They are periodically added to the pool and never reset. League exploiters also use a traditional self-play method PFSP to play against all policy pool agents, added to the pool if they show a high win rate and potentially reset to expose global blind spots. Main exploiters only compete with main agents to improve their robustness, are added to the pool after achieving a high win rate or specific training steps, and are reset upon each addition. Among those three agent types, the main agent is the core agent and embodies the final AlphaStar strategy. However, the introduction of three agent types significantly increases the training computation. Further studies (Han et al., 2020; Wang et al., 2021; Huang et al., 2024) have enhanced the league self-play training procedure.
4.3.2. MOBA Games
@@
OpenAI Five (Berner et al., 2019) defeated the world champion team in a simplified version of Dota 2, which supported only 17 heroes and excluded complex features like summons and illusions. The training process introduces a self-play method that combines two traditional self-play algorithms: with an 80% probability of engaging in vanilla self-play and a 20% probability of employing a technique similar to the PFSP used in AlphaStar (Vinyals et al., 2019). This technique selects each policy from the policy pool based on its quality score, which is continuously updated from competition results throughout training. Higher quality scores increase the likelihood of a policy being selected. OpenAI Five also requires extensive training resources.
Another notable MOBA game, especially viral in China, is Honor of Kings. The 1v1 mode is conquered by Ye et al. (2020b), which boasts a significant winning rate against top professional players. It utilizes a traditional self-play method $\delta$-uniform self-play to train the RL networks. The 5v5 mode was later mastered by Ye et al. (2020a) as well. Unlike OpenAI Five, which supported only 17 heroes, this work expands the hero pool to 40 heroes, substantially increasing the number of possible lineup combinations. It proposes a new self-play method referred to as curriculum self-play learning (CSPL). Specifically, the training process is divided into three stages. The first stage involves training fixed lineups through self-play and utilizing human data to balance the two teams to aid policy improvements. The second stage employs multi-teacher policy distillation to produce a distilled model. The final stage uses this distilled model as the initial model for another round of self-play with randomly picked lineups. This approach defeats professional player teams. The self-play generated data is also used to learn compelling lineup drafting by utilizing MCTS and neural networks (Chen et al., 2021).
4.3.3. Google Research Football
@@
WeKick (Ziyang Li, 2020) wins the GRF competition on Kaggle (Google, 2020), which simplifies the game dynamics by allowing competitors to control only one player, either the ball carrier on offense or the nearest defender on defense. It employed self-play strategies similar to those used in league training (Vinyals et al., 2019). It initializes its opponent policy pool using strategies developed through RL and Generative Adversarial Imitation Learning (GAIL) (Ho and Ermon, 2016) to facilitate the training process.
@@
Further research delves into the full football game rather than the simplified version. Team-PSRO (McAleer et al., 2023) extends PSRO series of self-play algorithms to team games, outperforming baselines in the 4v4 version of the full GRF. In the context of the 11v11 version, where the goalkeeper is rule-based controlled, TiKick (Huang et al., 2021a) utilizes vanilla self-play to collect samples and then employs imitation learning. Fictitious Cross-Play (FXP) (Xu et al., 2023) proposes a new self-play method similar to AlphaStar (Vinyals et al., 2019). It introduces two populations: the main population and the counter population. Policies in the counter population improve solely by cross-playing with policies in the main population as opponents, while policies in the main population engage in playing with policies from both populations. FXP achieves a win rate of over 94% against TiKick. TiZero (Lin et al., 2023), a follow-up to TiKick, combines curriculum learning with FP and PFSP (Vinyals et al., 2019) to avoid expert data reliance, achieving a higher TrueSkill rating (Herbrich et al., 2006) than TiKick.
Table 2. Overview of representative studies in empirical analysis. Category Game Number of Perfect Number of Algorithms Self-play Method Search Expert Data Players Information Game States Board Game Chess 2 $\checkmark$ $10^{45}$ AlphaZero Vanilla SP $\checkmark$ $\times$ MuZero Vanilla SP $\checkmark$ $\times$ Go 2 $\checkmark$ $10^{360}$ AlphaGo FP $\checkmark$ $\checkmark$ AlphaGo Zero Vanilla SP $\checkmark$ $\times$ AlphaZero Vanilla SP $\checkmark$ $\times$ MuZero Vanilla SP $\checkmark$ $\times$ Stratego 2 $\times$ $10^{535}$ DeepNash R-NaD $\times$ $\times$ Card Game HULHE 2 $\times$ $10^{17}$ Cepheus CFR+ $\checkmark$ $\times$ NFSP NFSP $\times$ $\times$ HUNL 2 $\times$ $10^{164}$ DeepStack CFR+ $\checkmark$ $\times$ Libratus MCCFR $\checkmark$ $\times$ ReBel CFR-D / FP $\checkmark$ $\times$ Six-player No-limit Texas Hold’em 6 $\times$ $10^{164}$ Pluribus MCCFR $\checkmark$ $\times$ DouDiZhu 3 $\times$ $10^{76}\sim 10^{106}$ DeltaDou Vanilla SP $\checkmark$ $\times$ DouZero Vanilla SP $\times$ $\checkmark$ PerfectDou Vanilla SP $\times$ $\times$ Mahjong 4 $\times$ $10^{169}$ Suphx Vanilla SP $\checkmark$ $\checkmark$ Video Game StarCraft II 2 $\times$ / AlphaStar FP & PFSP $\times$ $\checkmark$ Dota 2 10 $\times$ OpenAI Five Vanilla SP & PFSP $\times$ $\times$ Honor of Kings 10 $\times$ MOBA AI $\delta$-uniform SP $\checkmark$ $\checkmark$ Google Research Football 22 $\checkmark$ TiZero FP & PFSP $\times$ $\times$
5. Open Problems and Future Work
@@
Self-play methods have shown remarkable performance by leveraging iterative learning and adapting to complex environments. However, significant challenges and open questions remain, presenting opportunities for further research.
5.1. Theoretical Foundation
@@
Although NE has been shown to exist in games with finite players and finite actions (Nash et al., 1950), computing NE with self-play algorithms in larger games remains challenging and consequently, many studies aim to achieve approximate NE (Li et al., 2024). However, in some cases, even computing an approximate NE is difficult (Daskalakis, 2013). Some research has resorted to alternative solution concepts, such as CE (Marris et al., 2021) and $\alpha$-rank (Muller et al., 2020). While many successful self-play algorithms are grounded in theoretical intuition, formal game-theoretic guarantees for their effectiveness in large, imperfect-information games remain limited. For instance, approaches such as AlphaGo (Silver et al., 2016), AlphaStar (Samvelyan et al., 2019), and OpenAI Five (Berner et al., 2019) achieve empirical success. However, under realistic constraints such as finite computational resources and limited exploration, their convergence or optimality remains difficult to formally characterize within standard game-theoretic frameworks. Future work may benefit from further narrowing the gap between empirical success and theoretical understanding, either by extending existing theoretical tools to more complex settings or by designing new self-play methods with provable guarantees under realistic assumptions.
5.2. Non-stationarity of the Environment
@@
5.3. Scalability and Training Efficiency
@@
The scalability of self-play methods faces significant challenges as the number of teams and players within those teams increases. As the number of participants grows, the complexity of interactions explodes. For example, in OpenAI Five (Berner et al., 2019), the hero pool size is limited to only 17 heroes. MOBA AI (Ye et al., 2020a) extends this to a 40-hero pool with the help of curriculum learning, but it still cannot cover the entire hero pool available in the actual game. One potential solution is leveraging players’ inherent connections to optimize the learning process. For instance, using graph-based models to represent and exploit the relationships between players can help manage and reduce the complexity of large-scale multi-agent environments. These scalability issues are fundamentally rooted in the limited training efficiency of self-play methods from two aspects including computation and storage. The first issue is computational efficiency, which is induced by the iterative nature of self-play where agents repeatedly play against themselves or past versions. Moreover, although forming more complex populations and competitive mechanisms (Samvelyan et al., 2019) can enhance the intensity and quality of training, it further exacerbates the demand for computational resources. Techniques such as parallel computing, distributed learning, and more efficient neural network architectures could be explored to address these challenges. The second issue is storage because self-play requires maintaining a policy pool. Even when using a shared network architecture, storing the parameters of large models can be problematic. This issue is particularly pronounced in regret-minimization-based self-play algorithms, which must store the regrets for each information set and potential action. Managing both the computational load and storage requirements is essential for improving the training efficiency and scalability.
5.4. With Large Language Models
@@
With their remarkable capability and emergent generalizability, large language models (LLMs) have been regarded as a potential foundation for achieving human-level intelligence (Achiam et al., 2023), and self-play methods have been proposed to fine-tune LLMs, enhance LLMs’ reasoning performance, and build LLM-based agents with strong decision-making abilities. Post-training fine-tuning (Ouyang et al., 2022; Bai et al., 2022) is a key step in aligning LLM with more desired behaviors, but it requires a huge amount of human-annotated data. To reduce reliance on human-annotated data, Self-play fIne-tuNing (SPIN) (Chen et al., 2024) introduces a self-play mechanism to generate training data using the LLM itself and fine-tuning the LLM by distinguishing self-generated responses from human-annotated data. Another line of work (Munos et al., 2024; Swamy et al., 2024; Wu et al., 2024) formulate the alignment problem as a two-player constant-sum game and use self-play methods to solve the game. Some other work (Huang et al., 2023; Yuan et al., 2024) also utilizes model-generated data to fine-tune LLMs with minimal human annotations. The idea of self-improvement has also been applied to improve the reasoning ability of LLMs. Self-Play of Adversarial Game (SPAG) (Cheng et al., 2024) observes that self-play in a two-player adversarial language game called Adversarial Taboo can boost the LLM’s performance on various reasoning benchmarks. Besides improving the capability of LLMs, self-play methods have also contributed to building LLM-based agents with strong strategic abilities. A representative work is Cicero (Meta et al., 2022), which achieves human-level play in the game of Diplomacy by combining language models with an RL policy trained by self-play. Cicero uses the self-play policy to produce an intended action and prompts the language model to generate languages conditioned on the policy’s intent. Another line of work (Xu et al., 2024, 2025a) also combines LLM with self-play to train strategic language agents in the Werewolf game using reinforcement learning or counterfactual regret minimization. Despite recent progress, applying self-play with LLMs is still underexplored and requires further research.
5.5. Real-World Applications